OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (anonymous):

Dot product! \[ \vec u \cdot \vec v = | \vec u ||\vec v | \cos \theta\]

OpenStudy (anonymous):

so we know that u dot v / |u||v| = cos(t)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Sorry, I take this class online, it's really hard for me to understand what im doing...

OpenStudy (anonymous):

No worries :) The dot product is an operation that multiplies the x terms and the y terms (and z, etc...) of two vectors and adds them together. \[ \vec u \cdot \vec v = (2)(3) + (-4)(-8)\]

OpenStudy (anonymous):

It's also called the "scalar product" because it takes two vectors and generates a scalar number from their components.

OpenStudy (anonymous):

The dot product is also a measure of the angle between two vectors, so, like before \[ \vec u \cdot \vec v = |\vec u||\vec v| cos \theta \]

OpenStudy (anonymous):

so 6 + 32 = 38?

OpenStudy (anonymous):

is that what that means?

OpenStudy (anonymous):

Yeah. That's the left hand side. For the right hand side you take the magnitude of the two vectors and multiply them together. \[ |\vec u| = \sqrt{2^2 + (-4)^2} \\ |\vec v| = \sqrt{3^2 + (-8)^2}\] Then that goes in front of the cos

OpenStudy (anonymous):

Okayyyyy...

OpenStudy (anonymous):

Have you not seen any of the before ?

OpenStudy (anonymous):

Yea, I have, it's just always been confusing for me because of me not having a teacher to explain it to me. Literally all I have in the way of help is a bunch of words, like a text book but not as good...

OpenStudy (anonymous):

As long as it's not all a bunch of new gibberish :) And it's confusing no matter what, so don't feel bad. What about the stuff above makes sense? Or at least is less confusing?

OpenStudy (anonymous):

It's less confusing yea, so thanks! I can probably figure out the rest now, I hope lol

OpenStudy (anonymous):

Well I can help if you need clarification. Just holler :)